POWER SCALING OF VUV AND XUV FREQUENCY COMBS

\underline{DYLAN YOST}, ARMAN CINGOZ, JUN YE, JILA, National Institute of Standards and Technology and University of Colorado, Department of Physics, University of Colorado, Boulder, CO 80309-0440; AXEL RUEHL, MARTIN E. FERMANN, INGMAR HARTL, IMRA America, Inc., 1044 Woodridge Ave., Ann Arbor, MI 48105, USA.

By coupling a train of low-phase-noise, ultrashort pulses to a femtosecond enhancement cavity, it is possible to obtain a large enhancement of the pulse energy and subsequently drive the high harmonic generation process at ≥100 MHz repetition rates. At these repetition frequencies, the generated VUV-XUV frequency combs could potentially be used for a multitude of experiments requiring short wavelength radiation with exceptional temporal coherence and/or high average photon flux. We will present recent progress, including power scaling of the driving field with a corresponding factor of 10 improvement in harmonic power over previous results, as well as initial coherence tests of the generated VUV frequency comb.